Suppose borrow $100,000 at a rate of 4% (annual) with a 30 year maturity. Suppose rate...

  1. Suppose borrow $100,000 at a rate of 4% (annual) with a 30 year maturity. Suppose rate is locked for 5 years, and will reset to 6% at year 6. What will be the monthly payment at year 6 when the loan resets?

    a.   $477.18

    b. $612.02

    c. $582.72

    d. $1040.02

7.7 points   

QUESTION 5

  1. 5.   Suppose borrow $100,000 at a rate of 8% with a 30 year fixed mortgage. Suppose after 5 years the rate goes down to 5.5% and you’re considering refinancing it to a new 30 year fixed loan at the new rate.. Suppose the bank will charge a 2% pre payment penalty on the existing balance and a $1,500 fee on the new loan.   Suppose I will take the fees and add them into my new loan. What is my new monthly payment after the “refi”?

    a.   $733.76

    b. $1019.33

    c. $424.18

    d. $558.80

7.7 points   

QUESTION 6

  1. 6.   Buy a house for $200,000 with an 80% LTV. Suppose after one year the house goes up by 4%. Suppose the following year the house goes up by another 3%, and suppose he last year the house goes down by 2%. What’s my cash on cash return?

    a.   13.52%    b. 20%              c. 7.68%              d. -14.3%               e. 5.93%
Answer #1

4)

4 Loan due in 5 years 5 Loan due in 5 years 6 PV*(1+r)^n-P*((1+r)^n-1))/ $ 90,447.51 | = (100000*(1+(4%/12))^60)-PMT(4%/12,30

Monthly payment [Pxrx(1+r)^n] =((1+r)^n-1] Here, 1 Interest rate per annum 2 Number of years 3 Number of compoundings per per

*Please rate thumbs up